OpenStudy (anonymous):

Write the equation of a line in point-slope form that is parallel to the line y = 3x +6 and passes through the point (-5, 8)

OpenStudy (anonymous):

Point-slope form is an line equatin in the form of\[y – y _{1} = m(x – x_{1})\]Where: y,x - generic x and y thet you and up with in the end y1, x1 - coordinates of a point m - slope of the line so you have: x1 = -5 y1 = 8 m = 3 (from y = 3x + 6, cos it's parallel the slope is the same) so the equation is: y-8 = 3(x+5)